# Benchmarking Content-Based Puzzle Solvers on Corrupted Jigsaw Puzzles

### v1

@@ 4.3 Experiment 2: Corrupting Jigsaw Puzzles with Eroded Edges
When jigsaw puzzles are corrupted with eroded edges, the performance of all three heuristics and both deep learning base models declines rapidly, as shown in Fig. 3. Among these, Yu et al.’s solver achieves the highest accuracies, followed by Paikin and Tal’s and Gallagher’s heuristic. The Positional Diffusion and Transformer base models are on par with Gallagher on the Direct Comparison Metric, but perform significantly worse on the other metric. However, as the level of corruption increases to 30% or more, the differences between the approaches are diminishing and all five methods fail to reconstruct most puzzles.
Fine-tuning the deep learning models on this type of corruption again significantly improves their performance. Notably, Positional Diffusion becomes the best-performing method, outperforming both Yu et al.’s solver and the Transformer model in nearly all scenarios. However, it still faces challenges on larger puzzles, successfully reconstructing fewer than half of them when the edges are corrupted. The visual example of Tab. 2 confirms the previous findings. There, Gallagher’s solver fails to reassemble a quadratic image while the other methods misplace multiple pieces that are severely affected by this type of corruption. The prediction by Positional Diffusion is almost perfect.
@@ 4.4 Experiment 3: Corrupting Jigsaw Puzzles with Eroded Contents
Figure 4 presents the performance of the jigsaw puzzle solvers across different levels of content erosion, controlled by $\mathcal{E}$, for all four puzzle sizes. Among the heuristics, Gallagher’s solver performs the worst, followed by Paikin and Tal’s method, with Yu et al.’s heuristic achieving the highest accuracies. However, the performance for all three heuristics declines rapidly and they struggle to reassemble most test images when faced with eroded contents. While the base models of the deep learning approaches follow a similar downward trend, their Direct Comparison Metric remains higher for smaller puzzles up to $8\times 8$ pieces.
Fine-tuning the deep learning models on this type of corruption significantly improves their performance. Positional Diffusion remains the most effective solver, successfully reconstructing the highest number of puzzles in all scenarios. The Transformer model, while less effective than the diffusion model, still outperforms all other solvers by a considerable margin. A visual example is presented in Tab. 2 where the heuristics misplace every piece. The reconstructed image by the Transformer model shows more consistent structures while also misplacing many pieces. Positional Diffusion almost perfectly assembles the original image.
Table 2: Visual examples of the predictions from the three heuristics and two fine-tuned deep learning solvers on three types of corruptions. Each example is a $12\times 12$ puzzle from the PuzzleWikiArts dataset [14]. Pieces placed at the wrong absolute position are marked with red dots.

### v2

@@ 4.3 Experiment 2: Corrupting Jigsaw Puzzles with Eroded Edges
When jigsaw puzzles are corrupted with eroded edges, the performance of all three heuristics and both deep learning base models declines rapidly, as shown in Fig.˜3. Among these, Yu et al.’s solver achieves the highest accuracies, followed by Paikin and Tal’s and Gallagher’s heuristic. The Positional Diffusion and Transformer base models are on par with Gallagher on the Direct Comparison Metric, but perform significantly worse on the other metric. However, as the level of corruption increases to 30% or more, the differences between the approaches are diminishing and all five methods fail to reconstruct most puzzles.
Fine-tuning the deep learning models on this type of corruption again significantly improves their performance. Notably, Positional Diffusion becomes the best-performing method, outperforming both Yu et al.’s solver and the Transformer model in nearly all scenarios. However, it still faces challenges on larger puzzles, successfully reconstructing fewer than half of them when the edges are corrupted. The visual example of Tab.˜2 confirms the previous findings. There, Gallagher’s solver fails to reassemble a quadratic image while the other methods misplace multiple pieces that are severely affected by this type of corruption. The prediction by Positional Diffusion is almost perfect.
@@ 4.4 Experiment 3: Corrupting Jigsaw Puzzles with Eroded Contents
Figure˜4 presents the performance of the jigsaw puzzle solvers across different levels of content erosion, controlled by $\mathcal{E}$, for all four puzzle sizes. Among the heuristics, Gallagher’s solver performs the worst, followed by Paikin and Tal’s method, with Yu et al.’s heuristic achieving the highest accuracies. However, the performance for all three heuristics declines rapidly and they struggle to reassemble most test images when faced with eroded contents. While the base models of the deep learning approaches follow a similar downward trend, their Direct Comparison Metric remains higher for smaller puzzles up to $8\times 8$ pieces.
Fine-tuning the deep learning models on this type of corruption significantly improves their performance. Positional Diffusion remains the most effective solver, successfully reconstructing the highest number of puzzles in all scenarios. The Transformer model, while less effective than the diffusion model, still outperforms all other solvers by a considerable margin. A visual example is presented in Tab.˜2 where the heuristics misplace every piece. The reconstructed image by the Transformer model shows more consistent structures while also misplacing many pieces. Positional Diffusion almost perfectly assembles the original image.
Table 2: Visual examples of the predictions from the three heuristics and two fine-tuned deep learning solvers on three types of corruptions. Each example is a $12\times 12$ puzzle from the PuzzleWikiArts dataset [talon_ganzzle_2022]. Pieces placed at the wrong absolute position are marked with red dots.